What is an equation of the line that passes through the point (-3,-11) and is parallel to the line whose equation is 2x-y=4?
1
Expert's answer
2012-01-06T08:08:24-0500
Let's find the slope of the first line:
2x - y = 4, or y = 2x - 4, so k = 2.
Now we need to find the line with slope k=2 and passing througt the point (-3,-11):
y = kx+b
Let's substitute the point (-3,-11):
-11 = 2*(-3) + b ==> b = -5
So, y - 2x = -5 is the equation of the required line.